OpenStudy (mathstudent55):

Given y=sin(x^2 + y). Find dy/dx by implicit differentiation.

OpenStudy (mathstudent55):

\(y = \sin(x^2 + y) \) \(y' = (2x + y')\cos(x^2 + y)\) \(y' = 2x\cos(x^2+y)+y'\cos(x^2+y)\)

myininaya (myininaya):

Looks good so far.

OpenStudy (mathstudent55):

\(y' - y'\cos(x^2+y) = 2x\cos(x^2 +y)\) \(y'[1 -\cos(x^2+y)] = 2x\cos(x^2 +y)\)

OpenStudy (mathstudent55):

\(y' = \dfrac{ 2x\cos(x^2 +y) }{ 1 -\cos(x^2+y) }\)

myininaya (myininaya):

1-cos^2(u) can be written as sin^2(u) if you wanted

OpenStudy (anonymous):

Do you need the work checked?

OpenStudy (mathstudent55):

Great, thanks.

OpenStudy (carson889):

It is correct to the best of my knowledge

myininaya (myininaya):

also sin^2(x^2+y) on bottom

myininaya (myininaya):

and if you wanted you could write sin^2(x^2+y) as y^2 since y=sin(x^2+y) just if you wanted

myininaya (myininaya):

\[1-\cos^2(x^2+y)=\sin^2(x^2+y) =y^2\]

myininaya (myininaya):

oh wait there is no square on that cosine on bottom don't listen to me

OpenStudy (mathstudent55):

Here we had 1 - cos u in the denom, not 1 - cos^2 u, so it's not sin^2 u in the denom, right?

OpenStudy (asnaseer):

@myininaya - I don't see a \(\cos^2\) term in the original answer that mathstudent55 wrote down which was:\[y' = \dfrac{ 2x\cos(x^2 +y) }{ 1 -\cos(x^2+y) }\]

myininaya (myininaya):

yes yes yes i said don't listen to me already

myininaya (myininaya):

lol

OpenStudy (mathstudent55):

I'll leave the answer as this, then: \(y' = \dfrac{ 2x\cos(x^2 +y) }{ 1 -\cos(x^2+y) }\)

OpenStudy (asnaseer):

:)

myininaya (myininaya):

yes mathstudent is fine as is

OpenStudy (mathstudent55):

Thanks to everyone for your help!

OpenStudy (anonymous):

How did you get 1-cos(x^2 + y) in the denom? I get stuck on that part.

myininaya (myininaya):

rjaza are you good with \[y' - y'\cos(x^2+y) = 2x\cos(x^2 +y) \]

myininaya (myininaya):

i mean do you understand up until that point?

OpenStudy (anonymous):

Yes I got that part I subtracted dy/dx cos(x^2 +y) from both sides

OpenStudy (anonymous):

just don't understand where 1-cos comes from ?

myininaya (myininaya):

ok so if you factor the y' out on the left hand side then you have \[y'(1-\cos(x^2+y))=2x \cos(x^2+y)\]

myininaya (myininaya):

to solve that for y' you need to divide both sides by (1-cos(x^2+y))

OpenStudy (anonymous):

how does the -1 happen?

myininaya (myininaya):

-1?

OpenStudy (anonymous):

I did factor out dy/dx

OpenStudy (anonymous):

yes 1-cos(x^2 + y) where does the 1-cos come from. Sorry not -1

myininaya (myininaya):

\[\text{ factoring example } x-x^2=x(1-x)\]

myininaya (myininaya):

Both of these terms in the example have a common factor x

OpenStudy (anonymous):

Oh okay I see now!

myininaya (myininaya):

coolness

OpenStudy (anonymous):

thank you :)

myininaya (myininaya):

np